OpenStudy (anonymous):

Gerald can complete a math assignment in 2 hours and a physics assignment in 1 hour. If he has 8 hours to complete a set of assignments, how many math assignments can he complete if he decides to leave the science assignments for the next day? This situation is given by 2x + y = 8, where x is the number of completed math assignments and y is the number of completed science assignments. (x 0, y 0, x and y take only integer values) A) 1 B) 2 C) 3 D)4 E) 5 @iGreen @Kemist since ... yall are my favs XD

OpenStudy (anonymous):

You know: x is the number of completed math assignments y is the number of completed physics assignments AND 2x + y = 8 So, if "he decides to leave the physics assignments for the next day" how many science assignments will he do? Put that number in for y. After that you solve for x.

OpenStudy (anonymous):

@Kemist you make things so difficult ._. ....

OpenStudy (anonymous):

im stupid ........ you know ..... el- stupido ... .-.

OpenStudy (anonymous):

1 Math assignment = 2 hours. 4 Math assignments = 8 hours. 2(4)+y = 8

OpenStudy (igreen):

This is easy..you don't need an equation.. If he leaves out the physics assignments, and only does math assignments. And math assignments = 2 hours. He has 8 hours to do assignemnts. 8 / 2 = 4 math assignments.

OpenStudy (anonymous):

D ._. @SkylarLee

OpenStudy (anonymous):

oooooooooooooooh yall im just like under pressure i have 15 minutes to answer 40 more questions ... ;-;

OpenStudy (anonymous):

thank you @Kemist thank you @iGreen

OpenStudy (anonymous):

You're welcome, now I need help :D

OpenStudy (anonymous):

lol does it invole makeup ?? if not im clueless jk ......